How many moles of KBrO3 are required to prepare 0.0700 moles of Br2 according to the reaction: KBrO3 + 5KBr + 6HNO3 6KNO3 + 3Br2 + 3H2O *



A. 0.210


B. 0.0732


C. 0.0704


D. 0.220

1
Expert's answer
2022-04-11T17:02:06-0400

n(KBrO3)/n(Br2) = 1/3

n(KBrO3) = n(Br2)/3

n(KBrO3) = 0.0700 moles / 3 = 0.0233 moles